OpenStudy (anonymous):

Which of the following shows that polynomials are closed under subtraction when polynomial 5x − 6 is subtracted from 3x^2 − 6x + 2?

OpenStudy (anonymous):

Is somebody gonna answer my question?

OpenStudy (anonymous):

Nobody likes polynomials? wassup? I need help

OpenStudy (ivettef365):

you have 3x^2-6x+2-(5x-6) which is the same as 3x^2-6x+2-5x+6 now combine like terms

OpenStudy (anonymous):

your the best i bet everybody asks you for help huh?

OpenStudy (anonymous):

uh 3x+6x+5x = 14x and 6+2+2= 10

OpenStudy (ivettef365):

3x^2 can not be combined because you do NOT have another number with x^2 now -6x-5x can be combined because both have only ONE "X" so you have -11x and of course you can combine +2+6 = 8 both are just digits with no variables so you have for answer 3x^2-11x+8 Do you understand how to combine like terms ???

OpenStudy (anonymous):

yea i get it now thank you again